s=0.5(70-25)40=
can u solve this for me and explain it please

Answer by AnlytcPhil(1810)

s=0.5(70-25)40=
can u solve this for me and explain it please

   s = 0.5(70-25)40

1. Do whatever there is to do inside the first parentheses you come to going
   left to right. Copy everything else over.

2. If any parentheses remain with anything to do inside them, repeat step 1.
   If there are no parentheses which have anything to do inside them, do the
   first exponent, going left to right. Copy everything else over.

3. If any exponents remaimj, repeat step 2. If there are no exponents, do the
   first multiplication or division, going left to right. Copy everything else
   over.

4. If any multiplications or divisions remain, repeat step 3. If there are no
   multiplications, do the first addition or subtraction, going left to right.
   Copy everything else over.

5. If any additions or subtractions remain, repeat step 4. If there are no
   additions or subtractions you are done.

s = 0.5(70-25)40

1. The first (only) parentheses that has something to do in it is (70-25),
   so we replace (70-25) by (45), and copy everything else over:

s = 0.5(45)40
   
2. There are no exponents, so we go to 3

3. The first multiplication or division is 0.5(45) so we replace 0.5(45)
   by (22.5), and copy the rest, the 40, over.

s = (22.5)40

There is still a multiplication, so we repeat step 3, and get

s = 900

There is nothing left to do, so that's the answer.

Edwin
AnlytcPhil@aol.com